OpenStudy (anonymous):

David made an error when determining the product of (–2.5)(–8.3). Step 1: (–2.5)(–8.3) = (–8.3)(–2.5) Step 2: = (–8.3)(–2) + (–8.3)(–0.5) Step 3: = (–16.6) + (–4.15) Step 4: = –20.75 In which step did David make his first error?

OpenStudy (mathmath333):

\(\large -1\times -1=+1\)

OpenStudy (dan815):

step 3 ^ for that reason

OpenStudy (dan815):

-8.3 = -1 * 8.3 -2 = -1*2 -8.3 * -2 = -1*-1 *8.3*2 = 1*16.6 = 16.6